# The Shaped Transformer: Attention Models in the Infinite Depth-and-Width Limit

### v1

@@ 1 Introduction
Scaling limits have been used successfully to provide guidance on architecture [16, 17, 18] and tuning hyperparameters settings [19]. Our work represents a contribution in this direction. The ability to use such limits to diagnose instabilities depends on their tractability and faithfulness to real-world (finite) networks. In this regard, not all limits are created equal. In particular, the faithfulness of scaling limits depends critically on how other parameters are scaled with width. One of the simplest (and thus most popular) limits to work with – the “NTK” limit [20, 21, 22, 23, 24] – treats the depth of the network as fixed. As a result, at initialization, this limit does not accumulate sufficient random fluctuations over the depth of the network, leading to deterministic covariance matrices that do not agree with those of standard (finite) networks. Such networks have another defect: they are incapable of learning features in the limit [25]. Various other limits have been studied, towards identifying tractable yet faithful models of initialization and/or training. These include mean field limits [26, 27, 28, 29] and the perturbative regime [30, 31, 32, 33, 34].
This work operates in a relatively new regime – the proportional infinite depth-and-width limit – where depth $d$ and width $n$ diverge as the ratio $d/n$ tends to a positive constant. This limit, first studied by [35, ], has been the recent subject of intense study [35, 36, 37, 38, 39, 18]. This regime retains the network’s stochasticity and, at initialization, has been shown to closely resemble the behaviour of finite architectures, yet still yield a relatively simple limiting description, expressible in terms of stochastic differential equations [38, 18]. In this work, we fully characterize the initial output distributions of a network with skip connections and Softmax-based attention mechanisms, in the proportional infinite-depth-and-width limit.
@@
    We propose shaped attention, where we modify the Softmax-based attention mechanism to be a perturbation of the identity. We demonstrate that shaped attention successfully prevents the degeneracy of correlation in contrast to existing Transformer architectures (Figure 1).
@@ 2 Background
In deep learning theory, researchers have long sought to understand how networks internally represent different inputs and how different architectural choices affect these representations. The approach followed by work on signal propagation has been to study how the relative alignment of different inputs evolves across the network, as measured by the neural covariance $V^{\alpha\beta}_{\ell}:=\frac{1}{n}\langle x^{\alpha}_{\ell},x^{\beta}_{\ell}\rangle$ (or $\rho^{\alpha\beta}:=(V^{\alpha\alpha}_{\ell}V^{\beta\beta}_{\ell})^{-1/2}V^{\alpha\beta}_{\ell}$ if interested only in the correlation). At initialization, characterizations of this covariance structure have been exploited to infer important properties of neural networks [10, 11]. As an example, in the sequential infinite-width-then-depth limit, the correlation $\rho^{\alpha\beta}_{d}$ of MLPs is known to converge to a fixed point independent of the input [11, 16]. In this regime, the model is not able to discriminate different data points, which severely hinders training, as the gradient step for the deep layers is taken in the same direction regardless of the input. In the context of Softmax-based attention models, [8, ] proved that the feature matrix $X_{\ell}$ loses rank doubly exponentially fast with depth, and [9, ] showed how this leads to vanishing gradients of the queries and keys parameters, thus further highlighting how the stability of forward and backward passes are deeply entangled.
@@ 4.2 Shaped Attention
In Figure 3, we show how removing any of the proposed changes individually alters the neural covariance structure, which becomes degenerate for large depths, while the proposed modifications remains stable.
@@ 4.3 Main Result – Neural Covariance SDEs for Shaped Attention Models and Shaped Transformers
$b(V_{t}^{\alpha\beta})=\frac{\gamma^{2}}{\tau_{0}^{2}}\left[\frac{1}{m^{2}}\sum_{\nu,\kappa=1}^{m}V^{\nu\kappa}S_{1}^{\alpha\nu,\beta\kappa}+\frac{1}{2m}\sum_{\nu=1}^{m}(V^{\beta\nu}S_{2}^{\alpha\nu}+V^{\alpha\nu}S_{2}^{\beta\nu})\right]_{\alpha\leq\beta}\,,$
@@ 4.4 On Finite Time Stability of the SDE and Shaped Attention Networks
Although we did not observe numerical instability in majority of our simulations of the shaped attention networks and the corresponding SDE, we did observe that the drift component $b(V_{t})$ in Theorem 4.2 is cubic in the entries of $V_{t}$. Whenever the drift is not Lipschitz as in this case, we do not have general guarantees for the existence of a solution for all time (see the Feller test for explosions [52, Theorem 5.5.29]). In fact, MLPs with smooth activations also yield non-Lipschitz drift coefficients as seen in [18, ].
However, locally Lipschitz coefficients are sufficient to guarantee the existence of local solutions, in the sense of up to a stopping time [53, Proposition 6.9]. Not only does this fact help us establish a precise notion of convergence (Definition 4.1), we can also study the practical implications of this for finite sized attention networks. More specifically, we can inspect the effect of architectural changes to a stopping time.
@@ Appendix A Preliminaries: Covariance SDE Framework
Firstly, we will define the Skorohod $J_{1}$-topology, or just the Skorohod topology for short [62, Appendix 5]. The topology is used to describe convergence of continuous time processes with discontinuities, in particular, Markov chains with a continuous time interpolation fits in this category. Let $S$ be a complete separable metric space, and $D_{\mathbb{R}_{+},S}$ be the space of càdlàg functions (right continuous with left limits) from $\mathbb{R}_{+}\to S$. We use $x_{n}\xrightarrow{ul}x$ to denote locally uniform convergence (uniform on compact subsets of $\mathbb{R}_{+}$), and consider the class of bijections $\lambda$ on $\mathbb{R}_{+}$ so that $\lambda$ is strictly increasing with $\lambda_{0}=0$ (can be interpreted as a time change).
@@
In particular, we call the topology that induces this convergence the Skorohod topology [62, Theorem A5.3].
@@
Then it was shown that in the limit as $n,d\to\infty$ with $\frac{d}{n}=T>0$, the covariance matrix $V_{\ell}=\frac{1}{n}X_{\ell}X_{\ell}^{\top}$ (or equivalent the post activation $\frac{c}{n}\sigma_{s}(X_{\ell})\sigma_{s}(X_{\ell})^{\top}$ since the activation becomes infinitesimal) converges the solution of the following SDE in the Skorohod topology [18, Theorem 3.2]
@@ Appendix B SDE for Residual Network
For the cross product terms, we get (removing the dependence on $\ell$):
@@ C.2 Shaping the Softmax
The first few terms of the expansion provides a good approximation of $A_{\ell}$ when $\tau$ is large. Observe that for fixed $m$, the term $\frac{1}{m}11^{\top}$ is not of $\mathcal{O}(n^{-1})$. In particular, we can show that:
$\gamma^{2}\sum_{\delta,\omega=1}^{m}\mathbb{E}_{\ell}[A^{\alpha\delta}A^{\beta\omega}]V^{\delta\omega}=\gamma^{2}V^{\alpha\beta}+\frac{\gamma^{2}}{m}\sum_{\omega}V^{\alpha\omega}+\frac{\gamma^{2}}{m}\sum_{\delta}V^{\delta\beta}+\frac{\gamma^{2}}{m^{2}}\sum_{\delta\omega}V^{\delta\omega}+\mathcal{O}(\tau^{-1}).$
@@ C.4 Neural Covariance SDE for Stable Attention
Recall that:
@@ Appendix D Preliminary Experiments
We also incorporate the $\frac{1}{\sqrt{n}}$ factor into the initialization of the queries and keys weights by decreasing the variance of their entries by a factor of $\frac{1}{n}$. This allows us to not re-tune the learning rate for the queries and keys. We stress that at at initialization, the two formulations ($\tau=\sqrt{nn_{k}}$ and $\tau=\sqrt{n_{k}}$ with decreased weight’s variance) are equivalent. In both alternatives we train all the skip connection parameters $\lambda$, $\gamma$, and initialize $\gamma$ in the grid $(0.05,0.1,0.2)$ and set $\tau_{0}=1$. We report training instabilities (loss divergence) for larger values of $\gamma$. All models —including the baselines — use Adam [65] with learning rate warmup of $4000$ steps, the learning rate is tuned in the grid $(0.0001,0.0005,0.001,0.005)$. We report the train/test loss after $100K$ optimization steps, averaged over $2$ random seeds. All the other experimental details can be found in Section D.1.
@@
Learn 0.05 $2.26_{\pm 0.01}$ $2.10_{\pm 0.02}$
0.1 $2.21_{\pm 0.02}$ $2.08_{\pm 0.03}$ $2.21_{\pm 0.02}$ $2.02_{\pm 0.03}$
0.2 $2.20_{\pm 0.02}$ $2.07_{\pm 0.01}$ $2.28_{\pm 0.01}$ $2.09_{\pm 0.03}$
Recover 0.05 $2.22_{\pm 0.03}$ $2.04_{\pm 0.01}$
0.1 $\mathbf{2.19}_{\pm 0.04}$ $2.05_{\pm 0.02}$ $2.18_{\pm 0.02}$ $\mathbf{1.99}_{\pm 0.02}$
0.2 $\mathbf{2.19}_{\pm 0.03}$ $\mathbf{2.03}_{\pm 0.01}$ $2.17_{\pm 0.02}$ $2.00_{\pm 0.01}$
Baseline $2.26_{\pm 0.04}$ $2.19_{\pm 0.01}$ $\mathbf{2.16}_{\pm 0.02}$ $2.02_{\pm 0.04}$
Table 1: train/test loss of the proposed shaped attention, both in the "Recover" and "Learn" alternatives. Baseline refers to the vanilla Pre-LN Transformer. We report the best run under the learning rates $(0.0001,0.0005,0.001,0.005)$, averaged over $2$ random seeds. We do not perform the case $\gamma=0.05$ for the deeper network ($d=24$). We include the confidence interval of $\pm$ one standard deviation.
The Shaped Transformer Allows for Larger Learning Rates.
In Figure 5, we compare the training loss curves of standard Transformers with our shaped variants when varying the learning rate. For Softmax attention-based Transformers, we observe a degradation in training speed at $lr=0.0005$, and a divergent behavior at $lr=0.001$. On the other hand, shaped attention-based Transformers exhibit the same convergent behaviour across these learning rates. We observe the same phenomenon for both variants of shaped attention proposed here, and for all $\gamma\in[0.005,0.1,0.2]$. For completeness, we tried with even larger learning rates, observing that the network’s training diverges at $lr=0.005$.
(a) $d=18$
(b) $d=24$
Figure 5: Training loss curves across $3$ different learning rates for standard Transfomers (dashed lines) and our shaped attention in the recover setting (solid lines). The two plots indicate different depths (18 blocks on the left, 24 blocks on the right).

### v2

@@ 1 Introduction
Scaling limits have been used successfully to provide guidance on architecture [17, 18, 19] and tuning hyperparameters settings [20]. Our work represents a contribution in this direction. The ability to use such limits to diagnose instabilities depends on their tractability and faithfulness to real-world (finite) networks. In this regard, not all limits are created equal. In particular, the faithfulness of scaling limits depends critically on how other parameters are scaled with width. One of the simplest (and thus most popular) limits to work with – the “NTK” limit [21, 22, 23, 24, 25] – treats the depth of the network as fixed. As a result, at initialization, this limit does not accumulate sufficient random fluctuations over the depth of the network, leading to deterministic covariance matrices that do not agree with those of standard (finite) networks. Such networks have another defect: they are incapable of learning features in the limit [26]. Various other limits have been studied, towards identifying tractable yet faithful models of initialization and/or training. These include mean field limits [27, 28, 29, 30] and the perturbative regime [31, 32, 33, 34, 35, 36, 37].
This work operates in a relatively new regime – the proportional infinite depth-and-width limit – where depth $d$ and width $n$ diverge as the ratio $d/n$ tends to a positive constant. This limit, first analyzed by [38, ], has been the recent subject of study in the context of neural network [39, 40, 41, 42, 19]. A related line of work also studied the Lyapunov exponent for products of random matrices [43, 44, 45, 46]. This regime retains the network’s stochasticity and, at initialization, has been shown to closely resemble the behaviour of finite architectures, yet still yield a relatively simple limiting description, expressible in terms of stochastic differential equations [41, 19]. In this work, we fully characterize the initial output distributions of a network with skip connections and Softmax-based attention mechanisms, in the proportional infinite-depth-and-width limit.
@@
(a) Value Weights $W^{V}_{\ell}$
(b) Query Weights $W^{Q}_{\ell}$
Figure 2: Comparing gradients norms at initialization for different parameters as a function of depth, with and without shaped attention. The architecture is the same as in Figure 1 but with autoregressive causal masking, and the task is next-token prediction on code data. Left: Value weights $W^{V}_{\ell}$ for shaped attention, standard Pre-LN, and the original Post-LN block [2]. Right: the same gradient norm plot but for Query weights $W^{Q}_{l}$. We find that shaping the attention mechanism successfully prevents gradients from vanishing, while unshaped Transformers suffer from rapidly vanishing gradients. Interestingly, only the Post-LN query gradients vanish, but value gradients are stable across depths, which is consistent with the findings of [9, ]. On the other hand, shaped attention has stable gradients for both parameters inside and outside the Softmax nonlinearity.
@@
    We propose shaped attention, where we modify the Softmax-based attention mechanism to be a perturbation of the identity. We demonstrate that shaped attention successfully prevents the degeneracy of correlation in contrast to existing Transformer architectures (Figure 1). The enhanced stability in the forward pass is reflected in the gradients, which are also stable with depth, as we empirically show in Figure 2.
@@ 2 Background
In deep learning theory, researchers have long sought to understand how networks internally represent different inputs and how different architectural choices affect these representations. The approach followed by work on signal propagation has been to study how the relative alignment of different inputs evolves across the network, as measured by the neural covariance $V^{\alpha\beta}_{\ell}:=\frac{1}{n}\langle x^{\alpha}_{\ell},x^{\beta}_{\ell}\rangle$ (or $\rho^{\alpha\beta}:=(V^{\alpha\alpha}_{\ell}V^{\beta\beta}_{\ell})^{-1/2}V^{\alpha\beta}_{\ell}$ if interested only in the correlation). At initialization, characterizations of this covariance structure have been exploited to infer important properties of neural networks [10, 11]. As an example, in the sequential infinite-width-then-depth limit, the correlation $\rho^{\alpha\beta}_{d}$ of MLPs is known to converge to a fixed point independent of the input [11, 17]. In this regime, the model is not able to discriminate different data points, which severely hinders training, as the gradient step for the deep layers is taken in the same direction regardless of the input. In the context of Softmax-based attention models, [8, ] proved that the feature matrix $X_{\ell}$ loses rank doubly exponentially fast with depth, and [9, ] showed how this leads to vanishing gradients of the queries and keys parameters, thus further highlighting how the stability of forward and backward passes are deeply entangled (see also Figure 2).
@@ 4.2 Shaped Attention
In Figure 4, we show how removing any of the proposed changes individually alters the neural covariance structure, which becomes degenerate for large depths, while the proposed modifications remain stable. We stress that here for simplicity we focus on attention without masking. Shaped attention can be extended to include masking (e.g. casual masking) by centering each i-th row of the Softmax matrix by a different factor $1/m_{i}$, where $m_{i}$ is the number of un-masked tokens in the i-th row.
@@ 4.3 Main Result – Neural Covariance SDEs for Shaped Attention Models and Shaped Transformers
$b(V)=\frac{\gamma^{2}}{\tau_{0}^{2}}\left[\frac{1}{m^{2}}\sum_{\nu,\kappa=1}^{m}V^{\nu\kappa}S_{1}^{\alpha\nu,\beta\kappa}+\frac{1}{2m}\sum_{\nu=1}^{m}(V^{\beta\nu}S_{2}^{\alpha\nu}+V^{\alpha\nu}S_{2}^{\beta\nu})\right]_{\alpha\leq\beta}\,,$
@@ 4.4 On Finite Time Stability of the SDE and Shaped Attention Networks
Although we did not observe numerical instability in majority of our simulations of the shaped attention networks and the corresponding SDE, we did observe that the drift component $b(V_{t})$ in Theorem 4.2 is cubic in the entries of $V_{t}$. Whenever the drift is not Lipschitz as in this case, we do not have general guarantees for the existence of a solution for all time (see the Feller test for explosions [58, Theorem 5.5.29]). In fact, MLPs with smooth activations also yield non-Lipschitz drift coefficients as seen in [19, ].
However, locally Lipschitz coefficients are sufficient to guarantee the existence of local solutions, in the sense of up to a stopping time [59, Proposition 6.9]. Not only does this fact help us establish a precise notion of convergence (Definition 4.1), we can also study the practical implications of this for finite sized attention networks. More specifically, we can inspect the effect of architectural changes to a stopping time.
@@ Appendix A Preliminaries: Covariance SDE Framework
Firstly, we will define the Skorohod $J_{1}$-topology, or just the Skorohod topology for short [68, Appendix 5]. The topology is used to describe convergence of continuous time processes with discontinuities, in particular, Markov chains with a continuous time interpolation fits in this category. Let $S$ be a complete separable metric space, and $D_{\mathbb{R}_{+},S}$ be the space of càdlàg functions (right continuous with left limits) from $\mathbb{R}_{+}\to S$. We use $x_{n}\xrightarrow{ul}x$ to denote locally uniform convergence (uniform on compact subsets of $\mathbb{R}_{+}$), and consider the class of bijections $\lambda$ on $\mathbb{R}_{+}$ so that $\lambda$ is strictly increasing with $\lambda_{0}=0$ (can be interpreted as a time change).
@@
In particular, we call the topology that induces this convergence the Skorohod topology [68, Theorem A5.3].
@@
Then it was shown that in the limit as $n,d\to\infty$ with $\frac{d}{n}=T>0$, the covariance matrix $V_{\ell}=\frac{1}{n}X_{\ell}X_{\ell}^{\top}$ (or equivalent the post activation $\frac{c}{n}\sigma_{s}(X_{\ell})\sigma_{s}(X_{\ell})^{\top}$ since the activation becomes infinitesimal) converges the solution of the following SDE in the Skorohod topology [19, Theorem 3.2]
@@ Appendix B SDE for Residual Network
From now on, we will remove the dependence on $\ell$. Defining the residual branch as $f(X,W):=\frac{1}{\sqrt{n}}\sigma_{s}\left(\sqrt{\frac{c}{n}}XW^{\text{pre}}\right)W^{\text{post}}$, we can write $V^{\alpha\beta}$ as:
$V^{\alpha\beta}=\lambda^{2}X+\lambda\gamma\langle x^{\alpha},f(X,W)^{\beta}\rangle+\lambda\gamma\langle x^{\beta},f(X,W)^{\alpha}\rangle+\gamma^{2}\langle f(X,W)^{\alpha},f(X,W)^{\beta}\rangle.$
For the cross product terms, we get:
@@ C.2 Shaping the Softmax
Recall that we have the following model:
$X_{\ell+1}=\lambda X_{\ell}+\gamma A_{\ell}X_{\ell}\ \frac{1}{\sqrt{n}}W^{V}_{\ell}\,$
For the above model, $V_{\ell}$ has the following form:
@@
The first few terms of the expansion provides a good approximation of $A_{\ell}$ when $\tau$ is large. Observe that for fixed $m$, the zero order term $\frac{1}{m}11^{\top}$ is not of $\mathcal{O}(n^{-1})$. In particular, suppose that we use the attention model in Eq. 25. We can show that the expectation of the term in $\gamma^{2}$ (responsible for the drift) has the following form:
$\gamma^{2}\sum_{\delta,\omega=1}^{m}\mathbb{E}_{\ell}[A^{\alpha\delta}A^{\beta\omega}]V^{\delta\omega}=\gamma^{2}V^{\alpha\beta}+\frac{\gamma^{2}}{m}\sum_{\omega}V^{\alpha\omega}+\frac{\gamma^{2}}{m}\sum_{\delta}V^{\delta\beta}+\frac{\gamma^{2}}{m^{2}}\sum_{\delta\omega}V^{\delta\omega}+\mathcal{O}(\tau^{-1}),$
@@ C.4 Neural Covariance SDE for Stable Attention
Recall that we have the following model:
$X_{\ell+1}=\lambda X_{\ell}+\gamma A_{\ell}X_{\ell}\ \frac{1}{\sqrt{n}}W^{V}_{\ell}\,$
For the above model, $V_{\ell}$ has the following form:
@@ Appendix D Preliminary Experiments
We also incorporate the $\frac{1}{\sqrt{n}}$ factor into the initialization of the queries and keys weights by decreasing the variance of their entries by a factor of $\frac{1}{n}$. This allows us to not re-tune the learning rate for the queries and keys. We stress that at at initialization, the two formulations ($\tau=\sqrt{nn_{k}}$ and $\tau=\sqrt{n_{k}}$ with decreased weight’s variance) are equivalent. In both alternatives we train all the skip connection parameters $\lambda$, $\gamma$, and initialize $\gamma$ in the grid $(0.05,0.1,0.2)$ and set $\tau_{0}=1$. We report training instabilities (loss divergence) for larger values of $\gamma$. All models —including the baselines — use Adam [71] with learning rate warmup of $4000$ steps, the learning rate is tuned in the grid $(0.0001,0.0005,0.001,0.005)$. We report the train/test loss after $100K$ optimization steps, averaged over $4$ random seeds. All the other experimental details can be found in Section D.1.
@@
GLUE evaluation
Furthermore, we evaluate the trained models on three datasets from the GLUE benchmark [72] and summarize the results in Table 2. These demonstrate that our shaped Transformers holds promise by outperforming our pre-ln baseline.
@@
Learn 0.05 $2.09_{\pm 0.02}$ $2.07_{\pm 0.02}$ $2.03_{\pm 0.02}$ $2.03_{\pm 0.01}$
0.1 $2.10_{\pm 0.02}$ $2.07_{\pm 0.01}$ $\mathbf{2.02}_{\pm 0.01}$ $2.02_{\pm 0.03}$
0.2 $2.07_{\pm 0.05}$ $2.06_{\pm 0.02}$ $2.09_{\pm 0.04}$ $2.09_{\pm 0.03}$
Recover 0.05 $2.05_{\pm 0.03}$ $2.04_{\pm 0.02}$ $2.02_{\pm 0.04}$ $2.00_{\pm 0.01}$
0.1 $2.05_{\pm 0.01}$ $2.04_{\pm 0.01}$ $2.06_{\pm 0.03}$ $2.00_{\pm 0.02}$
0.2 $\mathbf{2.01}_{\pm 0.03}$ $\mathbf{2.03}_{\pm 0.01}$ $2.05_{\pm 0.04}$ $2.00_{\pm 0.01}$
Baseline $2.08_{\pm 0.03}$ $2.07_{\pm 0.01}$ $2.08_{\pm 0.03}$ $2.01_{\pm 0.01}$
Table 1: Training and test loss of the proposed shaped attention, both in the "Recover" and "Learn" alternatives. Baseline refers to the vanilla Pre-LN Transformer. We report the best run under the learning rates $(0.0001,0.0005,0.001,0.005)$, averaged over $4$ random seeds. We include the confidence interval of $\pm$ one standard deviation.
Model COLA MRPC RTE
$d=18$ Baseline $0.139_{\pm 0.008}$ $0.797_{\pm 0.010}$ $0.504_{\pm 0.024}$
Learn, $\gamma=0.2$ $0.182_{\pm 0.041}$ $0.813_{\pm 0.005}$ $0.528_{\pm 0.019}$
Recover, $\gamma=0.2$ $0.221_{\pm 0.042}$ $0.809_{\pm 0.009}$ $0.520_{\pm 0.029}$
$d=24$ Baseline $0.022_{\pm 0.055}$ $0.785_{\pm 0.009}$ $0.48_{\pm 0.046}$
Learn, $\gamma=0.05$ $0.150_{\pm 0.026}$ $0.812_{\pm 0.004}$ $0.513_{\pm 0.019}$
Recover, $\gamma=0.05$ $0.211_{\pm 0.039}$ $0.803_{\pm 0.012}$ $0.529_{\pm 0.019}$
BERT (Geiping et al.) 0.103 74.8 0.509
Table 2: Evaluation of the pretrained models on a subset of the GLUE benchmark. We take the checkpoints of the baselines and selected shaped Transformers ($\gamma=0.2$ for the shallower model $d=18$, $\gamma=0.05$ for the deeper model $d=24$). Then, we fine-tune on each of three datasets from GLUE (COLA, MRPC, RTE) for 5 epochs and Adam optimizer $lr=5e-4$ with batch size 16, and report test evaluation metric for the corresponding task.